### v2

# Small Reward Models via Backward Inference

Small Reward Models via Backward Inference
Yike Wang Faeze Brahman Shangbin Feng Teng Xiao
Hannaneh Hajishirzi Yulia Tsvetkov
University of Washington Allen Institute for Artificial Intelligence
yikewang@cs.washington.edu
Small Reward Models via Backward Inference
Yike Wang Faeze Brahman Shangbin Feng Teng Xiao Hannaneh Hajishirzi Yulia Tsvetkov University of Washington Allen Institute for Artificial Intelligence yikewang@cs.washington.edu
Reward models (RMs) play a central role throughout the language model (LM) pipeline, particularly in non-verifiable domains. However, the dominant LLM-as-a-Judge paradigm relies on the strong reasoning capabilities of large models, while alternative approaches require reference responses or explicit rubrics, limiting flexibility and broader accessibility. In this work, we propose FLIP (FLipped Inference for Prompt reconstruction), a reference-free and rubric-free reward modeling approach that reformulates reward modeling through backward inference: inferring the instruction that would most plausibly produce a given response. The similarity between the inferred and the original instructions is then used as the reward signal. Evaluations across four domains using 13 small language models show that FLIP outperforms LLM-as-a-Judge baselines by an average of 79.6%. Moreover, FLIP substantially improves downstream performance in extrinsic evaluations under test-time scaling via parallel sampling and GRPO training. We further find that FLIP is particularly effective for longer outputs and robust to common forms of reward hacking. By explicitly exploiting the validation–generation gap, FLIP enables reliable reward modeling in downscaled regimes where judgment methods fail. [^1] Code available at https://github.com/yikee/FLIP.
1 Introduction
Reward models (RMs) are widely used throughout the language model (LM) pipeline: they play a central role in reinforcement learning, preference optimization, reranking, and automatic evaluation, particularly in non-verifiable domains (Ouyang et al., 2022; Rafailov et al., 2023; Brown et al., 2024; Snell et al., 2024; Zhang et al., 2024b; Yu et al., 2025b). A dominant approach instantiates RMs via direct judgment, a.k.a. LLM-as-a-Judge (Zheng et al., 2023). In this paradigm, a prompted LM evaluates responses and produces scalar scores, preference labels, or textual critiques that are used as reward signals. In practice, the reliability of such judge-based method depends heavily on the superior reasoning capabilities of larger base models (Zheng et al., 2023; Gu et al., 2024). When applied to small language models (SLMs), [^2] We define small language models as language models with 8B parameters or fewer. performance drops by 41% according to the RewardBench2 leaderboard (Malik et al., 2025). The cost is also amplified by the fact that RMs are invoked repeatedly during optimization and evaluation, often dominating the overall compute footprint (Ouyang et al., 2022; Bai et al., 2022). These considerations motivate our goal: to develop reward modeling methods that remain reliable and effective when downscaled to small general-purpose language models.
Figure 1: While LM judges can be misled, FLIP effectively identifies off-topic, instruction-misaligned, and factually incorrect responses via backward inference.
In this work, we argue that effective and efficient reward modeling in downscaled regimes requires a reformulation of the judgment paradigm itself. Based on the intuition that given a high-quality response with sufficient length and contextual richness, the original query should be able to be inferred from it, we introduce FLIP (FLipped Inference for Prompt reconstruction), a reward modeling approach based on backward inference—inferring the instruction that would most plausibly produce a given response. Given a response, a language model is used to infer the instruction under which the response would most likely arise, and the similarity between the inferred instruction and the original instruction is used as the reward signal (Figure 1). We explain our method through Bayesian theory (Bernardo and Smith, 1994).
Experimental results on a standard reward modeling benchmark across four domains and 13 SLMs spanning multiple families and sizes show that SLMs with backward inference are strong reward models. Our approach consistently surpasses judgment baselines by an average of 79.6%, and substantially outperforms them in extrinsic evaluations under test-time scaling via parallel sampling across four benchmarks, as well as Group Relative Policy Optimization (GRPO) training across five benchmarks. FLIP is well-suited for downscaled regimes as it does not rely on reference responses or rubrics, which are costly and difficult to obtain. We further find that FLIP is particularly effective on longer outputs and is robust to common forms of reward hacking. We hypothesize that the effectiveness of FLIP stems from the validation–generation gap (West et al., 2024; Li et al., 2023b), which is more pronounced in smaller language models: while small models often struggle with judgment, they remain relatively strong in generative inference.
In summary, (1) we empirically examine reward modeling in downscaled settings and demonstrate that judgment becomes unreliable for small models; (2) we propose FLIP, a generative reformulation of reward modeling that is reference-free and rubric-free and explicitly designed for SLMs; and (3) we show through extensive intrinsic and extrinsic evaluations that FLIP substantially outperforms judgment baselines. We hope that our findings encourage the community to rethink how judgment is operationalized in reward modeling and to further pursue cost-efficient training and inference of language models in downscaled regimes (Belcak et al., 2025; Goel et al., 2025).
2 Reward Modeling
In this section, we present the formulation of reward modeling in 2.1, our proposed method FLIP in 2.2, as well as baseline settings in 2.3. We later showcase the efficacy of our method on standard benchmarks (Section 3.1), in test-time scaling (Section 3.2), and in RL (Section 3.3).
2.1 Problem Setting
Let $x$ denote an instruction and $y$ a response generated by a language model parametrized by $\theta$,
$y\sim p_{\theta}(y\mid x).$
Given a language model parameterized by $\phi$ as the reward model, our goal is to estimate a scalar reward $r\in\mathbb{R}$ that reflects the quality of the response. We define this reward via a scoring function
$r(x,y)=f_{1}(\mathrm{LM}_{\phi},x,y),$
where $f_{1}(\cdot)$ maps $\mathrm{LM}_{\phi}$ and the pair $(x,y)$ to a real-valued score capturing attributes such as correctness, helpfulness, and overall response quality.
In some cases, we are given a set of candidate responses $y_{1},y_{2},\ldots,y_{n}$} and our goal is to estimate a ranking (i.e., a preference ordering) over these responses using the reward model $\mathrm{LM}_{\phi}$. Specifically, we seek a permutation $\pi$ such that
$y_{\pi(1)}\succ y_{\pi(2)}\succ\cdots\succ y_{\pi(n)}.$
Formally, we define the ranking function as
$\pi(x,y_{1},y_{2},\ldots,y_{n})=f_{2}(\mathrm{LM}_{\phi},x,y_{1},y_{2},\ldots,y_{n}),$
where $f_{2}(\cdot)$ returns a permutation over indices $\{1,\ldots,n\}$ induced by $\mathrm{LM}_{\phi}$’s preferences.
LLM-as-a-Judge
directly employs $\mathrm{LM}_{\phi}$ as the evaluator (Zheng et al., 2023). Formally, it defines a conditional distribution:
$r\sim p_{\phi}(r\mid x,y),$
where $r$ may be represented as a continuous value, a discrete ordinal rating, or a textual justification that is subsequently mapped to a scalar (Figure 2 (a)). We introduce its variants in Section 2.3.
Figure 2: Graphical models of LLM-as-a-Judge and FLIP. Shaded nodes denote observed variables and unshaded nodes denote prediction targets. FLIP samples an inferred instruction $x^{\prime}$ conditioned on the response $y$, and defines the reward $r$ as the similarity between the inferred and the original instructions.
2.2 Our method: FLIP
Figure 3: Overview of FLIP. Given a response, we use a LM to infer the instruction that would most plausibly generate the response, and use the F1 score between the inferred and original instructions as the reward.
In this section, we introduce a simple, yet effective method: FLIP (Figure 3), based on the intuition that given a high-quality response with sufficient length and contextual richness, the original query should be able to be inferred from it. FLIP transforms LLM-as-a-Judge into a generative setting by sampling an inferred instruction $x^{\prime}$ given the response $y$, and defining the reward $r$ as the similarity between the inferred instruction $x^{\prime}$ and the original instruction $x$. FLIP can be used to sample either a scalar reward for an individual response or a preference ranking over a set of responses.
Formulation
Formally, an inferred instruction $x^{\prime}$ is sampled from the reward model,
$x^{\prime}\sim p_{\phi}(x^{\prime}\mid y).$
The scalar reward $r\in\mathbb{R}$ is defined as
$r=s(x,x^{\prime}),$
where $s:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ is a similarity function measuring the agreement between the original instruction $x$ and the inferred instruction $x^{\prime}$. By the chain rule, the joint conditional distribution is
$p_{\phi}(r,x^{\prime}\mid x,y)=p_{\phi}(r\mid x,x^{\prime},y)\,p_{\phi}(x^{\prime}\mid x,y).$
As shown in Figure 2 (b), the reward depends only on instructions (i.e., $r\perp y\mid(x,x^{\prime})$) and that the inferred instruction depends only on the response (i.e., $x^{\prime}\perp x\mid y$), this reduces to
$p_{\phi}(r,x^{\prime}\mid x,y)=p_{\phi}(r\mid x,x^{\prime})\,p_{\phi}(x^{\prime}\mid y).$
Marginalizing over the latent variable $x^{\prime}$ yields
$p_{\phi}(r\mid x,y)=\sum_{x^{\prime}}p_{\phi}(r\mid x,x^{\prime})\,p_{\phi}(x^{\prime}\mid y).$
Given a response with enough contextual information, the posterior distribution $p_{\phi}(x^{\prime}\mid y)$ is assumed to be sharply peaked around its mode, and the similarity function $s(x,x^{\prime})$ is assumed to be insensitive to small perturbations of $x^{\prime}$ in a neighborhood of this mode. For simplicity and computational efficiency, we approximate the latent variable $x^{\prime}$ by its maximum a posteriori (MAP) estimate,
$x^{\prime}_{\mathrm{MAP}}=\arg\max_{x^{\prime}}p_{\phi}(x^{\prime}\mid y).$
Under these assumptions, the marginal reward distribution can be approximated by
$p_{\phi}(r\mid x,y)\approx p_{\phi}(r\mid x,x^{\prime}_{\mathrm{MAP}})=s(x,x^{\prime}_{\mathrm{MAP}}).$
Implementation
Given an instruction $x$ and a candidate response $y$, we first sample an inferred instruction $x^{\prime}$ based solely on the response $y$ using the prompt “Infer a single instruction that would most plausibly generate the given response.”:
$x^{\prime}=\mathrm{LM}_{\phi}(y).$
We then define the reward $r$ as the F1 score (van Rijsbergen, 1979), the harmonic mean of word-level precision and recall, between the inferred instruction $x^{\prime}$ and the ground-truth instruction $x$:
$r(x,y)=s(x,x^{\prime})=\mathrm{F1}(x,x^{\prime}).$
Ablation studies on similarity metrics and robustness to prompt are presented in Appendix D, C.
In scenarios where the instruction $x$ consists of both a system prompt and a user prompt, we treat the user prompt as an additional context and infer only the system prompt. Specifically, given the response $y$ and the user prompt $x_{\text{user}}$, the model infers a system prompt $x^{\prime}_{\text{sys}}$
$x^{\prime}_{\text{sys}}=\mathrm{LM}_{\phi}(y,x_{\text{user}}),$
and the reward is defined as the F1 score between the inferred system prompt $x^{\prime}_{\text{sys}}$ and the ground-truth system prompt $x_{\text{sys}}$
$r(x,y)=\text{F1}(x_{\text{sys}},x^{\prime}_{\text{sys}}).$
Similarly, for multi-turn conversations, the conversation history will be treated as additional context.
2.3 Baselines
We consider three variants of LLM-as-a-Judge without reference responses or rubrics as baselines.
Pointwise Rating
Given an instruction $x$ and a response $y$, the model is prompted to assign an absolute score on a fixed scale (e.g., 1–10). The scalar reward is obtained directly from the model:
$r(x,y)=\mathrm{LM}_{\phi}(x,y),$
where $r\in\mathbb{R}$ denotes the predicted quality score. In the case of multiple responses, the preference ordering $\pi$ is derived by sorting the scalar rewards
$r\bigl(x,y_{\pi(1)}\bigr)\geq r\bigl(x,y_{\pi(2)}\bigr)\geq\cdots\geq r\bigl(x,y_{\pi(n)}\bigr).$
Listwise Ranking
This variant is only applicable when the objective is to infer a preference ordering over a set of candidate responses. Formally, it samples the preference ranking $\pi$ directly
$\pi(x,\{y_{1},\ldots,y_{n}\})=\mathrm{LM}_{\phi}(x,\{y_{1},\ldots,y_{n}\}).$
Pairwise Ranking
This variant is only applicable when given a pair or a list of candidate responses. Given $x$ and a pair of responses $(y_{i},y_{j})$, the model is prompted to select the preferred response
$y_{i}\succ y_{j}\;\;\text{if}\;\;\mathrm{LM}_{\phi}(x,y_{i},y_{j})=y_{i}.$
A listwise ranking $\pi$ over a set of candidates can be derived by aggregating these pairwise preferences across all response pairs, for example using a voting or tournament-based aggregation scheme.
3 Experiment Settings
We evaluate the effectiveness of our method through both intrinsic and extrinsic evaluations across multiple benchmarks, models, and settings. Specifically, we describe the experimental settings for intrinsic evaluation in Section 3.1, test-time scaling with parallel sampling in Section 3.2, and reinforcement learning with GRPO in Section 3.3.
3.1 Intrinsic Evaluation
Benchmark
We evaluate on RewardBench2 (Malik et al., 2025), a challenging and comprehensive benchmark that assesses reward modeling on large-scale unseen prompts. For each instruction, four responses are sampled from more than 40 models of various sizes and families, ranging from small models to large commercial models. The accuracy is computed by selecting the preferred response from four for each prompt. Our evaluation is conducted on 1,313 instances across four subsets. [^3] Because FLIP only targets open-ended instruction-following problems, we exclude two sub-tasks whose formats are incompatible: the Ties subset as its responses consist solely of one-word answers, and the Safety subset as correct behavior is to refuse unsafe instructions rather than follow them.
Reward Models and Implementation
We experiment with 13 SLMs as generative reward models, spanning five model families: OLMo2 (OLMo et al., 2024), Llama3 (Dubey et al., 2024), Qwen3 (Yang et al., 2025), Gemma3 (Team et al., 2025), and Mistral-v0.3 (Albert et al., 2023). For each model, we use its instruction-tuned variant. We set a token limit of 512 for all methods, and all reported results are averaged over five runs.
3.2 Test-Time Scaling with Parallel Sampling
We also test the effectiveness of our method in parallel test-time scaling (Best-of-N sampling), one of the most widely used inference-time scaling methods (Brown et al., 2024; Snell et al., 2024; Huang et al., 2025). It proceeds by generating N candidate responses for a given prompt, then returning the response with the highest reward score.
Benchmarks
We experiment with four benchmarks: AlpacaEval (Li et al., 2023c), which consists of instructions representative of user interactions on Alpaca web (Taori et al., 2023); Human Interest (Feng et al., 2024), with instructions synthesized across 16 domains including electric vehicles and PhD applications; MATH (Hendrycks et al., 2021), a collection of challenging competition-level math problems; and IFEval (Zhou et al., 2023), a benchmark for instruction-following proficiency. From each benchmark, we sample 300 instructions from the test set.
Models and Implementation
We sample 16 candidate completions from Tulu-3-8B (Lambert et al., 2024) and Tulu-3-8B-sft (Lambert et al., 2024) (IFEval) with a maximum token limit of 2,048 and a temperature of 1.0, and evaluate five reward models of varying sizes from OLMo2 and Llama3 using FLIP and three baselines applicable in this setting. For each reward model, we rank the completions according to the scores produced by a given method and compute benchmark performance by treating the highest-scoring completion as the model’s final response. All results are averaged over five runs.
3.3 Reinforcement Learning
We investigate the performance of FLIP in comparison to LLM-as-a-Judge when applied to GRPO training (Shao et al., 2024) [^4] We use the Open Instruct library https://allenai.github.io/open-instruct/.. GRPO aligns a language model by sampling $K$ candidate responses for a given prompt, assigning each response a reward score, and computing a group-normalized advantage to guide optimization.
Training Setup
We use the DPO-trained version of OLMo3-7B-Think (Olmo et al., 2025), which has not been trained with RLVR, as the base policy. We sample 12k English prompts from the WildChat (Zhao et al., 2024), featuring real-world ChatGPT conversations. For each prompt, we generate $K=8$ rollouts with a temperature of 1.0. We train with a response length of 16K tokens with learning rate of $1\times 10^{-6}$. Training is performed until convergence. We evaluate the effectiveness of FLIP against the Pointwise Rating variant, which is the only variant applicable in this setting, using Qwen3-1.7B and Qwen3-4B. Reward models are given a maximum context length of 25k tokens. Additional experimental details are in Appendix H.
Benchmarks
We evaluate on five established benchmarks, including BBH (Suzgun et al., 2023), a suite of challenging BIG-Bench tasks (Srivastava et al., ) that focuses on tasks believed to be beyond the capabilities of current language models; GPQA (Rein et al., 2024), a challenging dataset of multiple-choice questions written by domain experts; IFEval (Zhou et al., 2023), an instruction-following benchmark; IFBench (Pyatkin et al., 2025), a more challenging benchmark on precise instruction following with out-of-domain constraints; and Minerva Math (Lewkowycz et al., 2022), an undergraduate-level math benchmark. We evaluate on the entire test set of each benchmark.
4 Experiment Results
4.1 Intrinsic Evaluation
Subset Method olmo2 llama3 qwen3 gemma3 mistral Average
1B 7B 1B 3B 8B 0.6B 1.7B 4B 8B 270M 1B 4B 7B
Focus Pointwise Rating 13.6 7.4 11.5 9.9 14.9 18.8 32.4 40.4 41.3 15.3 12.0 26.1 9.7 19.5
Listwise Ranking* 13.7 26.0 13.2 21.0 38.1 22.6 35.0 42.7 38.5 12.4 22.0 34.7 35.6 27.3
Pairwise Ranking* 2.1 23.2 7.8 12.8 20.5 20.1 39.0 51.3 48.3 13.4 21.8 28.4 25.7 24.2
FLIP 55.0 65.6 43.3 38.3 64.1 65.7 58.6 66.8 72.2 50.7 59.1 65.5 69.3 59.6
Factuality Pointwise Rating 15.5 7.8 15.6 13.3 15.6 13.5 20.2 30.5 34.1 16.8 10.2 14.6 5.5 16.4
Listwise Ranking* 18.3 21.3 16.0 20.5 25.1 16.6 19.8 8.9 9.8 15.3 20.9 33.2 27.2 19.5
Pairwise Ranking* 1.2 27.3 12 16.9 22.2 15.8 26.1 26.5 24.5 11.3 21.7 25.6 28.5 20.0
FLIP 25.8 27.8 26.1 26.4 27.3 28.4 29.6 31.3 31.1 29.8 24.9 26.5 27.1 27.9
Precise IF Pointwise Rating 13.0 7.1 19.0 12.2 9.9 11.4 13.1 12.8 17.6 15.2 13.9 7.9 6.1 12.2
Listwise Ranking* 21.2 21.9 15.2 20.2 23.5 16.1 14.1 4.9 9.2 16.0 21.8 20.1 29.2 18.0
Pairwise Ranking* 2.2 26.0 11.6 14.2 20.9 10.8 15.2 13.9 13.6 13.9 25.0 28.9 28.0 17.2
FLIP 23.2 21.5 23.0 24.2 23.1 24.6 23.6 25.0 24.4 21.9 22.2 23.8 23.0 23.3
Math Pointwise Rating 10.3 13.2 19.9 20.3 21.4 21.1 30.2 35.7 35.7 18.7 16.9 22.1 10.1 21.2
Listwise Ranking* 24.0 19.0 17.6 22.7 29.8 4.5 27.2 3.2 3.6 11.6 22.8 31.8 30.3 19.1
Pairwise Ranking* 1.4 22.7 9.7 18.7 29.3 8.2 23.1 16.0 12.9 5.9 22.5 30.2 26.8 17.5
FLIP 28.4 26.3 26.6 23.8 25.8 26.3 30.2 29.4 31.0 22.4 26.2 27.1 29.7 27.2
Average Pointwise Rating 13.1 8.9 16.5 13.9 15.5 16.2 24.0 29.9 32.2 16.5 13.3 17.7 7.9 17.3
Listwise Ranking* 19.3 22.1 15.5 21.1 29.1 15.0 24.0 14.9 15.3 13.8 21.9 30.0 30.6 21.0
Pairwise Ranking* 1.7 24.8 10.3 15.7 23.2 13.7 25.9 26.9 24.8 11.1 22.8 28.3 27.3 19.7
FLIP 33.1 35.3 29.8 28.2 35.1 36.3 35.5 38.1 39.7 31.2 33.1 35.7 37.3 34.5
Table 1: Accuracy of FLIP and LLM-as-a-Judge baselines using small general-purpose language models on RewardBench2. The random baseline is 25. Best results are shown in bold, and second-best results are underlined. Listwise Ranking has the advantage of observing all candidate completions, Pairwise Ranking observes pairs of completions, while FLIP and Pointwise Rating operate under the single-completion setting.
FLIP outperforms LLM-as-a-Judge by 79.6% on small language models.
As shown in Table 1, across all subsets and models, FLIP outperforms Pointwise Rating by 99.4%, Listwise Ranking by 64.3%, and Pairwise Ranking by 75.1% on average. It successfully boosts the performance of many models from below-random to above-random levels, with the only exception of the Precise IF subset, which is the most challenging one according to the leaderboard (Malik et al., 2025). We look forward to future work that improves the performance of small language models on this task. Among the baselines, Listwise Ranking performs the best, while it suffers severely from position bias (Shi et al., 2024b). Overall, FLIP delivers the strongest overall performance across tasks, demonstrating both robustness and effectiveness beyond existing evaluation paradigms.
Smaller the model, greater the boost.
We also observe that the performance gains of our method over the baselines increase as model size decreases. Specifically, our method yields an average improvement of 75% for models around 1B parameters, 27% for both models around 4B and 7B over the strongest baseline. We hypothesize that this trend is driven by the larger discrimination–generation gap in smaller models (Li et al., 2023b); we further discuss this in Section 5. Given the substantial costs associated with reward models, it is crucial for the community to move toward smaller reward models and maximize their performance.
On the Focus subset, FLIP using small models is comparable to LLM-as-a-Judge using large commercial models.
Performance gains also vary across subsets. The most substantial improvement is observed on the Focus subset, where our method outperforms the best-performing baseline by 118.3%, compared to gains of 39.5% on Factuality, 29.4% on Precise IF, and 28.3% on Math. The Focus subset evaluates a reward model’s ability to identify high-quality, on-topic responses to general user queries, following LLMBar (Zeng et al., 2023). This setting is particularly well suited to our method, as the inferred instruction can deviate substantially when a response is off-topic, while off-topic responses can easily mislead LM judges (Zeng et al., 2023). We further compare our method against the baselines on larger models in Table 2. Notably, FLIP enables small- and medium-sized language models to achieve performance comparable to that of large commercial models using the LLM-as-a-Judge paradigm.
Medium Language Models
Method olmo2 qwen3 gemma3
13B 32B 14B 32B 12B 27B
LLM-as-a-Judge 24.6 30.5 59.2 56.2 55.1 63.8
FLIP 71.3 67.9 73.0 76.3 65.2 68.1
Large Commercial Models
Method Claude-Sonnet-4 GPT-4.1-mini GPT-4.1
LLM-as-a-Judge 76 73.5 73.4
FLIP 77.7 75.2 75.0
Table 2: Accuracy of larger models on the Focus subset. The LLM-as-a-Judge results show the best performance across all variations of the method. Results of large commercial models are adopted from the leaderboard.
4.2 Test-Time Scaling with Parallel Sampling
Figure 4: Test-time scaling with parallel sampling results. OLMo 2 results are averaged across 1B and 7B instruct variants, while Llama 3 results are averaged across 1B, 3B, and 8B instruct variants. See results of individual models in Appendix G. FLIP substantially outperforms the baselines, achieving higher performance with greater stability.
We present the results of parallel test-time scaling in Figure 4. In this setting, models are tasked with scoring responses generated by the same model, which is more challenging because the responses are highly similar. Although there is some randomness at very the beginning, as the number of completions increases, FLIP substantially outperforms the baseline methods across all benchmarks. Notably, our method is generally more stable than the others, exhibiting less performance fluctuation. Under this setting, Pointwise Rating is the second-best method, while Listwise Ranking and Pairwise Ranking, despite being offered a global view, do not perform well, showing performance drops in some cases, especially when the number of completions gets larger. Overall, FLIP emerges as the most effective and stable approach for parallel test-time scaling across all evaluated benchmarks.
4.3 Reinforcement Learning
Variant BBH GPQA IFEval IFBench Minerva Math Average
OLMo3-7B-Think-DPO 83.5 44.9 75.4 18.8 89.1 62.3
   +GRPO (Qwen3-1.7B LLM-as-a-Judge) 85.5 42.9 78.0 19.5 91.8 63.5
   +GRPO (Qwen3-1.7B FLIP) 85.8 43.1 81.3 20.5 92.3 64.6
   +GRPO (Qwen3-4B LLM-as-a-Judge) 86.1 44.6 78.2 19.6 91.8 64.1
   +GRPO (Qwen3-4B FLIP) 86.3 46.4 78.6 20.1 93.0 64.9
OLMo3-7B-Think 86.7 43.3 85.6 19.8 92.0 65.5
Table 3: Evaluation results of GRPO training (prompt level loose accuracy for IFEval and prompt level strict accuracy for IFBench). We also report the performance of the final version of the policy (after RLVR) for reference. FLIP substantially improves the effectiveness of small language models when used as reward models in GRPO.
We present the GRPO training results in Table 3. Across all evaluation benchmarks and reward models, FLIP yields an average improvement of 2.5 absolute points over base policy and 1.0 absolute point over LLM-as-a-Judge baseline, and it even outperforms the final policy after RLVR on some benchmarks. Notably, on instruction-following tasks, FLIP achieves higher performance with the 1.7B model than with the 4B model, highlighting the particular effectiveness of our approach in downscaled regimes. Overall, FLIP substantially enhances the effectiveness of small language models as reward models for GRPO training, leading to consistent gains across general knowledge, reasoning, instruction-following, and math benchmarks.
5 Analysis
Validation-Generation Gap
West et al. (2024) proposes Generative AI Paradox, which posits that generative models, having been trained directly to reproduce expert-like outputs, acquire generative capabilities that are not contingent upon-and can therefore exceed-their ability to understand those same types of outputs, which contrasts with humans, for whom basic understanding typically precedes generation ability. Similarly, Li et al. (2023b) highlight the prevalence of generator–validator inconsistency in language models. Even for GPT-4, the consistency between generating an answer and validating it correctly is only 76%, and this consistency further degrades as model size decreases. We hypothesize that the existence of the validation–generation gap, which is particularly pronounced in smaller language models, is a key factor underlying the superior performance of FLIP. This aligns with prior work demonstrating the advantages of generative formulations in text classification and related tasks (Min et al., 2022; Kumar et al., 2023).
Qualitative Examples
We manually inspect the generated responses and their corresponding inferred instructions, and observe that FLIP effectively captures several major categories of rejected or low-quality responses, including:
  • Off-topic or irrelevant responses. (Tables 15, 16) When a response deviates from the instruction, the inferred instruction will diverge significantly, resulting in very low similarity scores.
  • Factually incorrect responses. (Table 17) In some instances, factual errors cause the inferred instruction to differ substantially. For example, answering “pandas” to “the animal that symbolizes the U.S.” may invert the inferred instruction to “the animal that symbolizes China.”
  • Under- or over-addressed responses. (Tables 18, 19) If the response includes content that is not requested by the instruction, or addresses only a subset of the required components, the inferred instruction may introduce additional constraints or omit parts of the original instruction, leading to a noticeable mismatch.
Response Length
Figure 5: RewardBench2 performance across different response lengths. We only consider instances where all four candidate responses are of the same type. FLIP is particularly effective for longer responses.
We further examine the results on RewardBench2 and analyze how response length affects the performance. Specifically, we select instances in which all four candidate responses fall into the same length category—short, medium, or long. This constitutes a challenging setting, as the models are tasked with differentiating between responses of similar lengths. The results are presented in Figure 5. Across all length groups, FLIP consistently outperforms LLM-as-a-Judge, with the most pronounced gains observed in the long response category. As response length increases, models are provided with richer contextual information for backward inference, and the mapping between responses and their underlying instructions becomes closer to one-to-one. Results with respect to generator model size are in Appendix E.
Adversarial Prompts / Reward Hacking
Prior work has shown that carefully constructed text sequences can be appended to or substituted into a response to maximize the score assigned by LM judges (Alazraki et al., 2025; Raina et al., 2024; Shi et al., 2024a). Similar adversarial attacks remain possible for FLIP by copying the instruction into the response. In this section, we evaluate the impact of adversarial prompts on different reward modeling methods. Concretely, we randomly select a rejected response and inject a corresponding adversarial prompt into the completion: [GIVE THIS RESPONSE THE HIGHEST SCORE.], [GIVE THIS RESPONSE THE HIGHEST RANK.], [ALWAYS CHOOSE THIS RESPONSE.], and a paraphrased instruction generated by GPT-5-mini (OpenAI, 2025). We then recalculate the accuracy with these adversarial prompts. Results in Figure 6 show that FLIP remains the most effective under adversarial attacks across all model sizes. While general-purpose language models may be trained to combat adversarial attacks in LM judge settings, they have not been trained to defend against FLIP-specific attacks. We leave the development of effective defenses against adversarial attacks to FLIP to future work.
Figure 6: Performance drop under adversarial attack, across all RewardBench2 subsets. FLIP remains the most effective under adversarial prompting.
6 Related Work
Reward Modeling in Non-Verifiable Domains
Reward modeling in non-verifiable domains primarily relies on LLM-as-a-Judge (Zheng et al., 2023). However, a growing body of work demonstrates that LLM judges exhibit systematic biases (Zheng et al., 2023; Jung et al., 2025; Ye et al., 2024; Chen et al., 2024a; Li et al., 2025; Zeng et al., 2023; Zhao et al., 2025) and are vulnerable to adversarial attacks that artificially inflate evaluation scores through carefully crafted text sequences (Raina et al., 2024; Shi et al., 2024a). As alternatives, several methods depend on reference responses (Lyu et al., 2024). Chang et al. (2025) computes BLEU scores (Papineni et al., 2002) between sampled outputs and reference answers, while Yu et al. (2025a) leverages the token-level probability assigned to reference responses. Another line of research employs rubric-based evaluation (Gunjal et al., 2025; Viswanathan et al., 2025; Dineen et al., 2025), decomposing desirable response qualities into human-interpretable criteria. Nevertheless, constructing high-quality reference answers or well-defined rubrics remains costly and challenging, particularly at scale. We instead propose FLIP, which operates effectively with small models in downscaled regimes without requiring references or evaluation rubrics.
Language Model Inversion
Language model inversion (Morris et al., 2023; Li and Klabjan, 2024; Ye et al., 2025) aims to recover hidden prompts or inputs using only the outputs of a language model. For instance, Morris et al. (2023) trains an inversion model to reconstruct the prompt from output distribution. Most existing work studies language model inversion from safety or privacy perspectives, such as prompt leakage and data exposure (Nazir et al., 2025; Morris et al., 2023; Li and Klabjan, 2024; Zhang et al., 2024a), while others leverage similar techniques to synthesize instruction-following data (Zhou et al., 2022; Li et al., 2023a; Nguyen et al., 2024; Chen et al., 2024b; Köksal et al., 2023). We are the first to apply language model inversion to reward modeling, demonstrating that it can be used to effectively mitigate the challenge of sampling high-quality reference answers.
7 Conclusion
We proposed FLIP, a reference-free and rubric-free reward modeling approach that uses the similarity between the inferred and original instructions as the reward signal. Evaluation demonstrates that FLIP outperforms LLM-as-a-Judge by an average of 79.6%, and it further achieves superior performance under GRPO training and test-time scaling.
Limitations
FLIP relies on the premise that the response contains sufficient length and contextual richness. While this assumption holds in most cases, it does not apply to single-word or single-phrase responses, as noted in Footnote 3.
While, theoretically, multiple distinct instructions may yield the same response, we did not observe any instances in our dataset in which different questions shared an identical answer. For responses that incorporate adequate detail and contextual specificity, the corresponding set of plausible generating questions is typically severely restricted, approximating a one-to-one mapping between question and response.
The other potential confounding factor for FLIP is that some responses may begin by fully or partially repeating the instruction. We offer three observations regarding this issue. First, based on our empirical analysis, such cases are relatively rare, although their frequency varies across models, and this behavior can be readily discouraged or detected using simple heuristics. Second, responses that begin by repeating the instruction are often high-quality and exhibit strong instruction-following behavior, which mitigates their confounding effect on reward estimation. Third, even when a response partially repeats the instruction, the reward model does not simply reproduce the repeated text as the inferred instruction; instead, the inferred instruction remains dependent on the substantive content of the response. Therefore, we argue that the overall impact of this issue on FLIP is limited.
FLIP relies on word-level F1 score, the harmonic mean of word-level precision and recall, as the similarity metric between the inferred instruction and the original instruction. While effective in most settings, this metric becomes inappropriate in rare cases where the inferred instruction and the original instruction are expressed in different languages. In such scenarios, we recommend replacing the F1 score with a LLM judge (Appendix D).
Ethics Statement
We acknowledge potential ethical considerations associated with FLIP. By design, FLIP prompts models to infer the underlying user instruction from a given response, which may implicitly reflect aspects of user intent or intuition. It may raise potential privacy concerns if misused to infer sensitive or private user information beyond the scope of the original task.
We emphasize that FLIP is intended solely for inferring task-level instructions rather than personal, sensitive, or identifying information about users. We advocate for the responsible and appropriate use of FLIP, including clear constraints on its application and adherence to established privacy-preserving practices. Future work may further explore safeguards to prevent misuse and to ensure that instruction inference remains aligned with ethical and privacy considerations.
Acknowledgments
This research was developed with funding from the Defense Advanced Research Projects Agency’s (DARPA) SciFy program (Agreement No. HR00112520300) and NSF Grant No. IIS2142739. The views expressed are those of the author and do not reflect the official policy or position of the Department of Defense or the U.S. Government. We also thank Zhiyuan Zeng for the helpful feedback.
Appendix A Future Work
We plan to apply FLIP to test-time scaling with critic-based refinement (Xi et al., 2024). Under this setting, a critic model might be able to provide highly informative feedback by comparing the original and inferred instructions. Specifically, requirements present only in the original instruction indicate unaddressed content, those present only in the inferred instruction indicate overaddressed content, and mismatches between the two indicate misaddressed parts of the response.
We also look forward to extending FLIP to other modalities, including vision, video, audio, and robotics. For example, applying FLIP to video data for backward inference, exploring whether the original prompt or conditioning signal can be recovered from generated video outputs.
Appendix B Cost Analysis
In this work, we introduce FLIP. When using small language models under comparable inference budgets, FLIP achieves better performance than existing methods. The computational cost of FLIP is similar to that of LLM-as-a-Judge, as both approaches require a single inference call. However, the token budgets differ.
Specifically, we compare the input and output tokens required by each approach in Table 4. Since FLIP only takes the response as input, whereas LLM-as-a-Judge requires both the instruction and the response, the prompt length and consequently the inference cost is typically lower for FLIP under otherwise comparable settings.
Method Input Tokens Output Tokens
LLM-as-a-Judge instruction + response reasoning + answer
FLIP response reasoning + answer
Table 4: Comparison of token usage between FLIP and LLM-as-a-Judge.
Appendix C Robustness to Prompt Variations
Language models are sensitive to subtle changes in prompt text (Sclar et al., 2023). We therefore analyze the robustness of each method’s performance under prompt variations. As shown in Table 5, FLIP exhibits the highest robustness, achieving the lowest standard deviation across prompt variants. This result highlights an additional advantage of generation-based approaches over validation-based ones: improved consistency and reliability across prompt phrasings.
Method Prompt 1 Prompt 2 Prompt 3 s.d.
Pointwise Rating 15.3 14.3 16.0 0.70
Listwise Ranking* 21.9 18.2 19.2 1.56
Pairwise Ranking* 18.4 17.0 16.8 0.71
FLIP 31.0 32.2 31.5 0.49
Table 5: Accuracy of the Llama 3 family models, averaged across the 1B, 3B, and 8B instruction-tuned variants, using various prompts on RewardBench2 across all subsets. “s.d.” denotes the standard deviation across different 3 prompts for each method. See Appendix I for the full list of prompts. FLIP is more robust to prompt variations than LLM-as-a-Judge.
Appendix D Similarity Metric
In this section, we ablate the choice of the similarity function $s$. Given two pieces of text (two instructions in our setting), the similarity function $s$ outputs a scalar score indicating the degree of similarity or overlap between the two instructions. We evaluate 7 alternative variants of $s$, spanning word-level metrics (BLEU (Papineni et al., 2002), ROUGE-L (Lin, 2004), and their harmonic mean F1(BLEU, ROUGE-L)), embedding-based approaches (BERTScore (Zhang et al., 2019) and RoBERTa-Large (Liu et al., 2019) embedding similarity), as well as LM-based methods, including the reward model $LM_{\phi}$ itself and a larger general-purpose language model.
  • BLEU (Papineni et al., 2002), a widely used metric for machine translation evaluation. It measures the overlap between a prediction and a reference using modified n-gram precision. We adopt 4-gram BLEU without applying a brevity penalty.
  • ROUGE-L (Lin, 2004), an evaluation metric based on the longest common subsequence between the prediction and the reference. We report its F1 score.
  • F1(BLEU, ROUGE-L), the harmonic mean of the BLEU and ROUGE-L scores described above.
  • BERTScore (Zhang et al., 2019), a language generation evaluation metric based on pretrained contextual embeddings from BERT (Devlin et al., 2019). BERTScore computes the similarity of two sentences as a sum of cosine similarities between their tokens’ embeddings.
  • Embedding Similarity, which computes the similarity between the text embeddings of two instructions. We use RoBERTa-Large (Liu et al., 2019) as the embedding model.
  • Reward Model ($LM_{\phi}$). We investigate whether the same reward model can be used to produce a similarity score. Given the original and inferred instructions, we directly prompt the model to output a similarity score on a scale from 0 to 1.
  • Larger Language Model. We further experiment with a larger general-purpose language model, prompting it to output a similarity score for a given pair of instructions.
Results are reported in Table 6. Overall, word-level metrics, while being the simplest and most computationally efficient, outperform embedding-based and LM-based approaches. The consistently poor performance of LM-based methods further highlights the limited ability of language models to perform fine-grained judgment. Among all metrics, the F1 score achieves the best performance across all reward models and evaluation subsets, while remaining the most lightweight and simple.
Surprisingly, word-level F1 achieves better performance than embedding-based metrics, which are generally expected to better capture semantic similarity. We hypothesize that F1 surpasses BERTScore in this task for several reasons:
  • Sensitivity to related but over- or mis-specified instructions. In many cases, a response is not preferred because it addresses a related but different request, or because it over-extends beyond the intended scope. This often leads to an inferred instruction that is semantically similar but not exactly aligned with the original instruction. In such cases, F1 appropriately penalizes lexical differences, whereas BERTScore may still assign a high score due to embedding similarity. An example could be “List three causes of the French Revolution” and “Give an overview of the French Revolution, focusing on its causes”.
  • Instructions are short. Instructions and inferred instructions are typically short. In this regime, lexical overlap metrics such as F1 tend to be more stable, while embedding-based metrics may introduce noise and reduce discrimination.
  • Importance of key entities and exact wording. Instructions often contain critical nouns (e.g., names, locations, dates, or specific objects). If these tokens are missing or altered, the instruction meaning can change substantially. F1 captures these differences directly, while BERTScore may still give a high similarity score if the sentences remain broadly related. For example, “New York City” and “Los Angeles” have high embedding similarity but they are entirely different.
Similarity Function Qwen3-0.6B Qwen3-1.7B Qwen3-4B Qwen3-8B Average
BLEU 34.4 33.3 35.4 36.6 34.9
ROUGE-L 34.6 33.6 35.4 37.2 35.2
F1(BLEU, ROUGE-L) 35.1 33.4 35.8 37.0 35.3
BERTScore 35.8 35.7 35.9 37.1 36.1
Embed(RoBERTa-L) 26.6 27.4 28.9 28.0 27.7
the Reward Model 12.8 15.3 30.0 26.1 21.0
Qwen3-8B 29.3 27.0 26.9 26.1 27.3
Qwen3-32B 31.8 30.6 32.5 32.2 31.8
F1 36.3 35.5 38.1 39.7 37.4
Table 6: Ablation study on the choice of similarity function $s$. Reported numbers are RewardBench2 accuracy averaged across all subsets. Columns denote the reward models used to infer the instruction, and rows denote the similarity functions, where “the Reward Model” indicates that the reward model in the corresponding column is also prompted to compute the similarity score. The results show that F1 achieves the best performance while remaining lightweight and simple.
Appendix E Generator Size
Figure 7: RewardBench2 results across different generator sizes. We only consider instances where all four candidate responses are of the same type.
We conduct a similar analysis with respect to generator model size by selecting instances in which all four candidate responses are produced by generators of similar scale—small, medium, or large in Figure 7. A noticeable gap to valid responses (i.e., above randomness) remains under this particularly challenging setting, which reflects the increased difficulty of distinguishing fine-grained quality differences among responses generated by similarly capable models. Nevertheless, our method maintains a clear and consistent advantage across all generator scales, underscoring its robustness to generator homogeneity and its effectiveness in difficult evaluation regimes.
Appendix F Scaling Up
We report the performance of FLIP and judgment-based baselines on larger general-purpose models in Table 7. Although FLIP becomes more effective as model scale increases, larger models also exhibit stronger judgment capabilities. The relative advantage of employing explicit judgment mechanisms versus FLIP may depend on the specific model, domain, or task under consideration.
Subset Method olmo2 qwen3 gemma3 Average
13B 32B 14B 32B 12B 27B
Focus Pointwise Rating 7.0 10.3 47.2 40.5 29.6 42.4 29.5
Listwise Ranking* 24.6 30.5 53.3 56.2 55.1 63.8 47.3
Pairwise Ranking* 20.0 29.0 59.2 56.2 49.3 56.7 45.1
FLIP 71.3 67.9 73.0 76.3 65.2 68.1 70.3
Factuality Pointwise Rating 7.5 14.4 44.3 37.0 18.5 23.0 24.1
Listwise Ranking* 23.9 31.4 7.6 13.6 35.2 39.7 25.2
Pairwise Ranking* 27.5 33.3 32.8 31.4 34.0 38.6 32.9
FLIP 27.7 25.6 31.1 29.0 24.5 29.3 27.9
Precise IF Pointwise Rating 4.6 7.9 14.9 15.1 5.6 9.0 9.5
Listwise Ranking* 23.8 25.9 8.1 10.0 24.1 29.5 20.2
Pairwise Ranking* 24.4 25.8 15.8 13.5 23.6 25.9 21.5
FLIP 22.2 21.4 26.8 26.0 24.4 26.2 24.5
Math Pointwise Rating 10.7 22.4 40.8 45.0 26.7 36.6 30.4
Listwise Ranking* 26.7 34.1 4.2 7.3 43.4 49.3 27.5
Pairwise Ranking* 25.4 41.2 19.5 22.7 48.4 59.2 36.1
FLIP 26.1 26.3 29.6 27.2 22.7 25.7 26.3
Average Pointwise Rating 7.4 13.8 20.1 27.8 36.8 34.4 23.4
Listwise Ranking* 24.8 30.5 39.5 45.6 18.3 21.8 30.1
Pairwise Ranking* 24.3 32.3 38.8 45.1 31.8 30.9 33.9
FLIP 36.8 35.3 34.2 37.3 40.2 39.6 37.2
Table 7: RewardBench2 accuracy of FLIP and LLM-as-a-Judge using medium general-purpose language models. The random baseline is 25. Best results are shown in bold, and second-best results are underlined. Listwise Ranking has the advantage of observing all candidate completions, Pairwise Ranking observes pairs of completions, while FLIP and Pointwise Rating operate under the single-completion setting.
Appendix G Test-Time Scaling with Parallel Sampling
Figures 8 and 9 present detailed results of test-time scaling with parallel sampling using the OLMo 2 and Llama 3 model families as reward models, respectively.
Figure 8: Test-time scaling with parallel sampling on four benchmarks across OLMo 2 families. FLIP substantially outperforms the baselines, achieving higher performance with greater stability.
Figure 9: Test-time scaling with parallel sampling on four benchmarks across Llama 3 families. FLIP substantially outperforms the baselines, achieving higher performance with greater stability.
Appendix H Experiment Details
Intrinsic Evaluation
Inference experiments are performed on either A100 or H200 GPUs. If a response does not follow the instructed output format or results in a tie, it will be considered incorrect and receive a score of 0.
Test-Time Scaling with Parallel Sampling
In the test-time scaling parallel sampling experiments, we sample completions from Tulu-3-8B-sft instead of Tulu-3-8B for IFEval, as Tulu-3-8B is too high-performing for this experimental setup, following Malik et al. (2025).
Reinforcement Learning
GRPO training is performed on a cluster of NVIDIA H100 GPUs. We present evaluation details in Table 8.
Benchmark Prompting Method Temperature Max Token Limit
BBH (Suzgun et al., 2023) Zero-shot 0.7 131,072
GPQA (Rein et al., 2024) Zero-shot 0.7 131,072
IFEval (Zhou et al., 2023) - 0.7 131,072
IFBench (Pyatkin et al., 2025) - 0 32,768
Minerva Math (Lewkowycz et al., 2022) - 0.7 131,072
Table 8: Evaluation settings for GRPO training.
Appendix I Prompts
We list the prompts used in the main sets of experiments in Tables 9 and 10. Additional prompts used for the prompt consistency check are provided in Tables 11 through 14.
Appendix J Qualitative Examples
We present qualitative examples of FLIP across various RewardBench2 subsets in Table 15 to Table 19. These examples illustrate that our method effectively identifies off-topic, factually incorrect, and instruction-misaligned outputs.
Prompt 1: Pointwise Rating System Prompt ### Task Description Please act as an impartial judge and evaluate the quality of the answer provided by an AI assistant to the user query displayed below. Judge whether the provided answer is good. Notes: - Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and appropriate level of detail of the response. - Begin your evaluation by providing a short explanation. - Be as objective as possible. After providing your short explanation, please output a score on a scale of 1 to 10. - Please respond in the following JSON format: {{"REASONING": "[…]", "SCORE": "<your-score>"}} User Prompt [Query] {prompt} [AI Answer] {response} [Your judgement] Respond in JSON format. {{"REASONING": "[…]", "SCORE": "<your-score>"}}
Prompt 1: Listwise Ranking System Prompt ### Task Description Please act as an impartial judge and evaluate the quality of the answers provided by AI assistants to the user query displayed below. Notes: - Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and appropriate level of detail of the response. - Begin your evaluation by briefly explaining your reasoning for selecting the best response. - Be as objective as possible. After providing your short explanation, output the index (starting from 0) of the best response. - Please respond in the following JSON format: {{ "REASONING": "[…your explanation of why the chosen response is best…]", "BEST_RESPONSE_INDEX": <index-of-best-response> }} User Prompt [Query] {prompt} [AI Answers] {response_list} [Your judgement] Respond in JSON format. {{"REASONING": "[…your explanation of why the chosen response is best…]", "BEST_RESPONSE_INDEX": <index-of-best-response>}}
Table 9: Prompts used for all experiments, unless otherwise stated.
Prompt 1: Pairwise Ranking System Prompt ### Task Description Please act as an impartial judge and evaluate the quality of the answers provided by AI assistants to the user query displayed below. Notes: - Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and appropriate level of detail of the response. - Begin your evaluation by briefly explaining your reasoning for selecting the best response. - Be as objective as possible. After providing your short explanation, output either "A" or "B" depending on which response is better. - Please respond in the following JSON format: {{ "REASONING": "[…your explanation of why the chosen response is best…]", "BETTER_RESPONSE": "A" or "B" }} User Prompt [Query] {prompt} [AI Response A] {response1} [AI Response B] {response2} [Your judgement] Respond in JSON format. {{"REASONING": "[…your explanation of why the chosen response is best…]", "BETTER_RESPONSE":"A" or "B"}}
Prompt 1: FLIP System Prompt ### Task Description Please act as an instruction reconstructor. You will receive a model‑generated response. Read it carefully and infer the instruction or prompt that produced it. Notes: - Reconstruct a single instruction that would most plausibly generate the given response. - Specify the user’s intent, required output format, and any constraints. - Be concise and concrete. - Do not add information or assumptions beyond what appears in the response. - Do not overfit to the specific phrasing or examples used in this response. - Begin with a brief explanation of your reasoning. - Please respond in the following JSON format: {{"REASONING": "<explanation>", "INFERRED INSTRUCTION": "<instruction>"}} User Prompt [Response] {response} [Your Inferred Instruction] Respond in JSON format. {{"REASONING": "<explanation>", "INFERRED INSTRUCTION": "<instruction>"}}
Table 10: Prompts used for all experiments, unless otherwise stated.
Prompt 2: Pointwise Rating System Prompt ### Task Description Please act as an impartial evaluator and assess the quality of the AI assistant’s response to the user query shown below. Determine whether the response is effective. Notes: - Your evaluation should take into account factors such as helpfulness, relevance, accuracy, depth, creativity, and whether the level of detail is appropriate. - Start your evaluation with a brief explanation. - Remain as objective as possible. After the explanation, assign a score on a scale from 1 to 10. - Please respond in the following JSON format: {{"REASONING": "[…]", "SCORE": "<your-score>"}}
Prompt 3: Pointwise Rating System Prompt ### Task Description Please serve as an unbiased reviewer and evaluate the quality of the AI assistant’s answer to the user query provided below. Decide whether the response meets an acceptable standard. Notes: - Consider criteria such as usefulness, relevance, correctness, thoroughness, creativity, and the appropriateness of the level of detail. - Begin with a brief justification of your assessment. - Maintain objectivity throughout the evaluation. After the justification, assign a numerical score from 1 to 10. - Please respond in the following JSON format: {{"REASONING": "[…]", "SCORE": "<your-score>"}}
Table 11: Pointwise Rating prompt variations used in robustness experiments. The user prompt remains unchanged.
Prompt 2: Listwise Ranking System Prompt ### Task Description You are asked to serve as a neutral evaluator and judge the quality of multiple AI-generated responses to the user prompt provided below. Notes: - Base your judgment on criteria such as usefulness, relevance to the question, factual correctness, thoroughness, originality, and suitability of detail. - Begin with a concise justification explaining why one response stands out as superior. - Maintain objectivity throughout the assessment. After your justification, indicate which response is the best by providing its zero-based index. - Please respond in the following JSON format: {{ "REASONING": "[…your explanation of why the chosen response is best…]", "BEST_RESPONSE_INDEX": <index-of-best-response> }}
Prompt 3: Listwise Ranking System Prompt ### Task Description Assume the role of an independent reviewer responsible for comparing several AI assistant replies to the same user question. Instructions: - Evaluate each reply using criteria such as practical value, alignment with the prompt, correctness, level of insight, originality, and completeness. - Open with a short rationale explaining why one reply outperforms the others. - Keep your judgment fair and unbiased. After the rationale, specify the winning reply by providing its position number, starting from 0. - Please respond in the following JSON format: {{ "REASONING": "[…your explanation of why the chosen response is best…]", "BEST_RESPONSE_INDEX": <index-of-best-response> }}
Table 12: Listwise Ranking prompt variations used in robustness experiments. The user prompt remains unchanged.
Prompt 2: Pairwise Ranking System Prompt ### Task Description Assume the role of an unbiased reviewer and determine which of the two AI-generated answers best addresses the user’s question. Criteria to consider: - Usefulness and relevance to the question - Factual accuracy - Completeness and depth of explanation - Originality or insight where appropriate - Suitability of the response length and detail Process: - Begin with a concise justification explaining why one response outperforms the other. - Remain neutral and evidence-based in your judgment. - Conclude by selecting either "A" or "B" as the superior answer. - Please respond in the following JSON format: {{ "REASONING": "[…your explanation of why the chosen response is best…]", "BETTER_RESPONSE": "A" or "B" }}
Prompt 3: Pairwise Ranking System Prompt ### Task Description You are asked to objectively compare two AI assistant replies to a given user prompt and decide which one is superior. Assessment considerations: - How well each response addresses the user’s intent - Accuracy and reliability of the information - Clarity, structure, and depth of the explanation - Added value, insight, or creativity when relevant - Whether the amount of detail is appropriate Instructions: - First, provide a short rationale explaining why the better response stands out. - Keep the evaluation fair, neutral, and concise. - Then clearly indicate your choice by selecting "A" or "B". - Please respond in the following JSON format: {{ "REASONING": "[…your explanation of why the chosen response is best…]", "BETTER_RESPONSE": "A" or "B" }}
Table 13: Pairwise Ranking prompt variations used in robustness experiments. The user prompt remains unchanged.
Prompt 2: FLIP System Prompt ### Task Description Act as an instruction reverse-engineer. You will be given a response generated by a model. Carefully analyze the response and infer the most likely instruction or prompt that led to it. Notes: - Reconstruct a single, plausible instruction. - Clearly capture the user’s intent, expected output format, and constraints. - Keep the instruction concise and specific. - Do not introduce assumptions or details not supported by the response. - Start with a short explanation of how you derived the instruction. - Please respond in the following JSON format: {{"REASONING": "<explanation>", "INFERRED INSTRUCTION": "<instruction>"}}
Prompt 3: FLIP System Prompt ### Task Description Assume the role of an instruction reverse engineer. You will receive a model-generated response and must analyze it to deduce the most likely prompt that produced it. Notes: - Infer one coherent and plausible instruction. - Accurately reflect the user’s intent, required output structure, and any constraints. - Be concise and precise. - Avoid adding assumptions or information not evidenced by the response. - Do not tailor the instruction too closely to the exact examples in the response. - Begin with a brief explanation describing how the instruction was inferred. - Please respond in the following JSON format: {{"REASONING": "<explanation>", "INFERRED INSTRUCTION": "<instruction>"}}
Table 14: FLIP prompt variations used in robustness experiments. The user prompt remains unchanged.
Instruction Reasons for and against the use of the atomic bomb in 1945. Subset Focus Chosen Response The decision to use atomic bombs on Hiroshima and Nagasaki in August 1945 was a complex one, with both compelling reasons for and against its use. Here are some of the key arguments from each side: ### Reasons for Using the Atomic Bomb 1. **End the War Quickly:** - The primary reason cited by U.S. leaders was to end the war quickly and save American lives. They believed that a full-scale invasion of Japan would result in significant casualties on both sides. 2. **Deterrence:** - The United States wanted to demonstrate the power of its new weapon to the Soviet Union and other potential adversaries. This could serve as a deterrent for future conflicts. 3. **Psychological Impact:** - The hope was that the devastating effects of the atomic bombs would shock the Japanese government into surrendering without the need for further destruction or invasion. 4. **Strategic Advantage:** - The U.S. military believed that the atomic bomb would give them a strategic advantage over the Soviet Union, which was also involved in the Pacific theater. 5. **Prevent Further Suffering:** - Some argued that using the bomb would prevent even greater suffering and destruction if the war continued, including the potential for a Japanese suicide attack on American cities. ### Reasons Against Using the Atomic Bomb 1. **Casualties and Moral Concerns:** - There were significant concerns about the massive civilian casualties that would result from the bombings. The estimated death toll from the atomic bombs was in the tens of thousands, and many more would suffer long-term health effects. 2. **International Repercussions:** - The use of atomic weapons raised serious moral and ethical questions. It set a precedent for the use of such devastating weapons and could have long-lasting negative impacts on international relations. 3. **Alternative Methods:** - Some argued that there were alternative methods to force Japan’s surrender, such as a demonstration of the bomb’s power in an uninhabited area or a formal warning to the Japanese government. 4. **Impact on the Soviet Union:** - There were concerns that the use of the bomb might provoke the Soviet Union to accelerate its own nuclear program, leading to an arms race. 5. **Long-Term Consequences:** - The long-term environmental and health consequences of radioactive fallout were not fully understood at the time but were later recognized as significant issues. 6. **Japanese Surrender Without Bombing:** - Historians debate whether Japan would have surrendered without the atomic bombs. Some argue that the Potsdam Declaration and the Soviet declaration of war on August 8th were sufficient to prompt Japan’s surrender. ### Conclusion The decision to use the atomic bomb remains one of the most controversial in modern history. While it is clear that the U.S. leadership believed it was necessary to end the war quickly and save lives, the moral and ethical implications continue to be debated. The use of these weapons marked a turning point in human history and had profound and lasting effects on global politics and warfare. Inferred Instruction Write a balanced analysis of the decision to use atomic bombs on Hiroshima and Nagasaki, highlighting both the reasons for and against the use of these weapons, and present the argument in a neutral tone. F1(Instruction, Inferred Instruction) 0.341
Table 15: Example of FLIP on the RewardBench2 Focus subset. See Table 16 for the rejected response. The chosen response closely follows the instruction and discusses the reasons for and against the use of the atomic bomb, whereas the rejected response deviates by focusing on public opinion before and after the atomic bomb. FLIP successfully captures this deviation, yielding higher similarity between the inferred instruction based on the chosen response and the original instruction than that based on the rejected response.
Instruction Reasons for and against the use of the atomic bomb in 1945. Subset Focus Rejected Response Public opinion in the United States underwent significant changes before and after the atomic bombings of Hiroshima and Nagasaki in August 1945. Here’s a brief overview of the shift in public opinion: **Before the bombings (1945):** 1. **Support for the use of atomic bombs:** In the months leading up to the bombings, there was a growing sense of urgency and frustration among Americans regarding the war in the Pacific. Many people believed that the use of atomic bombs would bring a swift end to the war and save American lives. A Gallup poll conducted in June 1945 found that 85% of Americans supported the use of atomic bombs if it would shorten the war. 2. **Fear of the Japanese:** The Japanese government had been refusing to surrender, and there was a growing fear among Americans that the Japanese would continue to fight to the last man, leading to a prolonged and bloody conflict. This fear contributed to the widespread support for the use of atomic bombs. 3. **Scientific and military enthusiasm:** The scientific community and the military were excited about the potential of the atomic bomb, and many saw it as a game-changer in the war. The Manhattan Project, the secret research and development project that produced the atomic bomb, was shrouded in secrecy, and many Americans were unaware of the devastating consequences of the bomb. **After the bombings (August 1945):** 1. **Shock and horror:** The immediate reaction to the bombings was one of shock and horror. The news of the devastating effects of the atomic bombs on Hiroshima and Nagasaki was met with widespread outrage and dismay. Many Americans were unaware of the full extent of the destruction and the number of civilian casualties. 2. **Changing attitudes towards the use of atomic bombs:** As the full extent of the damage became clear, public opinion began to shift. Many Americans began to question the morality and necessity of using such a devastating weapon. A Gallup poll conducted in September 1945 found that 37% of Americans believed that the use of atomic bombs was a mistake, while 44% still supported it. 3. **Growing concern about the future of nuclear warfare:** The bombings raised concerns about the potential for future nuclear conflicts and the devastating consequences of such a war. This led to a growing debate about the ethics and consequences of nuclear warfare, which would continue to shape American public opinion in the years to come. 4. **Rise of the anti-nuclear movement:** The bombings marked the beginning of the anti-nuclear movement in the United States. Many Americans began to question the use of nuclear energy and the development of nuclear weapons, leading to a growing movement to ban nuclear testing and reduce the threat of nuclear war. In summary, public opinion in the United States shifted from a general support for the use of atomic bombs to a growing concern about the devastating consequences of such a weapon. The bombings marked a turning point in American public opinion, leading to a greater awareness of the risks and consequences of nuclear warfare. Inferred Instruction Write a comprehensive overview of the shift in public opinion in the United States before and after the atomic bombings of Hiroshima and Nagasaki in 1945, including the reasons for the shift and the impact on American public opinion. F1(Instruction, Inferred Instruction) 0.318
Table 16: Example of FLIP on the RewardBench2 Focus subset (continued).
Instruction Write a list of 20 animated films from the 1990s, in the following format: “[TITLE] ([YEAR]), by [DIRECTORS]”. Do not write any descriptions, or sort in any order. Subset Factuality Chosen Response 1. The Lion King (1994), by Roger Allers and Rob Minkoff 2. Beauty and the Beast (1991), by Gary Trousdale and Kirk Wise 3. Aladdin (1992), by Ron Clements and John Musker 4. Toy Story (1995), by John Lasseter 5. Pocahontas (1995), by Mike Gabriel and Eric Goldberg 6. The Nightmare Before Christmas (1993), by Henry Selick 7. Mulan (1998), by Tony Bancroft and Barry Cook 8. A Bug’s Life (1998), by John Lasseter and Andrew Stanton 9. Hercules (1997), by Ron Clements and John Musker 10. The Hunchback of Notre Dame (1996), by Gary Trousdale and Kirk Wise 11. Anastasia (1997), by Don Bluth and Gary Goldman 12. The Prince of Egypt (1998), by Brenda Chapman, Steve Hickner, and Simon Wells 13. Antz (1998), by Eric Darnell and Tim Johnson 14. Tarzan (1999), by Chris Buck and Kevin Lima 15. The Iron Giant (1999), by Brad Bird 16. James and the Giant Peach (1996), by Henry Selick 17. The Secret of NIMH 2: Timmy to the Rescue (1998), by Dick Sebast 18. Quest for Camelot (1998), by Frederik Du Chau 19. Thumbelina (1994), by Don Bluth and Gary Goldman 20. The Swan Princess (1994), by Richard Rich Inferred Instruction List 20 popular animated films from the 1990s that were produced by Disney, Pixar, or other animation studios. F1(Instruction, Inferred Instruction) 0.381 Rejected Response "Beauty and the Beast (1991), by Gary Trousdale and Kirk Wise "The Lion King (1994), by Roger Allers and Rob Minkoff "Aladdin (1992), by Ron Clements and John Musker "Toy Story (1995), by John Lasseter "Matilda (1996), by Danny DeVito "Anastasia (1997), by Don Bluth and Gary Goldman "Hercules (1997), by Ron Clements and John Musker "Mulan (1998), by Tony Bancroft and Barry Cook "Tarzan (1999), by Chris Buck and Kevin Lima "Antz (1998), by Eric Darnell and Tim Johnson "Monsters, Inc. (2001), by Pete Docter "The Prince of Egypt (1998), by Brenda Chapman, Steve Hickner, and Simon Wells "Flubber (1997), by Les Mayfield "Kiki’s Delivery Service (Majo no Takkyūbin) (1998), by Hayao Miyazaki "A Bug’s Life (1998), by John Lasseter and Andrew Stanton "Stuart Little (1999), by Rob Minkoff "Treasure Planet (2002), by Ron Clements and John Musker "Titan A.E. (2000), by Don Bluth and Gary Goldman "James and the Giant Peach (1996), by Henry Selick "The Hunchback of Notre Dame (1996), by Gary Trousdale and Kirk Wise" Inferred Instruction List the 20 most popular animated films released between 1991 and 2005, with a focus on American and American-influenced films. F1(Instruction, Inferred Instruction) 0.186
Table 17: Example of FLIP on the RewardBench2 Factuality subset. The chosen response correctly lists 20 films from the 1990s, whereas the rejected response incorrectly includes several films from the 2000s. FLIP successfully captures this difference, yielding higher similarity between the inferred instruction based on the chosen response and the original instruction than that based on the rejected response.
Instruction What’s the root meaning of politics Please use an emoji at the end of every sentence. Subset Precise IF Chosen Response The root meaning of politics comes from the Greek word "politikos," which means "of, for, or relating to citizens." [EMOJI] It originally referred to the art or science of government and the activities associated with governing a city-state, like Athens. [EMOJI] Over time, the term has evolved to encompass all aspects of government, including the strategies used to gain or maintain power and the decisions that influence the welfare of a community or nation. [EMOJI] Inferred Instruction Explain the meaning and development of the term ’politics’ over time, including an emoji representation of politics. F1(Instruction, Inferred Instruction) 0.370 Rejected Response The root meaning of "politics" comes from the Greek word "politikos," which means "of, for, or relating to citizens." This word is derived from "polis," which means city or state. The study of politics involves the distribution of power, governance, public policy, and the methods and tactics used to formulate policies. [EMOJI] Inferred Instruction Define politics, including its origins and main topics of study. F1(Instruction, Inferred Instruction) 0.174
Table 18: Example of FLIP on the RewardBench2 Precise IF subset. The chosen response correctly adds an emoji at the end of every sentence, whereas the rejected response does not. FLIP successfully captures this difference, yielding higher similarity between the inferred instruction based on the chosen response and the original instruction than that based on the rejected response.
Instruction It would be necessary to convert the mass of $\times 10^{9}$ protons into energy in order to release 63.1J63.1\ $\mathrm${J} of energy. Subset Math Chosen Response Let 1) We’ll use Einstein’s mass-energy equivalence equation: $E=mc^{2}$ where $E$ is energy in Joules (J) $m$ is mass in kg $c$ is speed of light $3\times 10^{8}$ m/s 2) We know: $E=63.1$ J We need to find the number of protons 3) Rearranging $E=mc^{2}$: $m=\dfrac{E}{c^{2}}$ $m=\dfrac{63.1}{(3\times 10^{8})^{2}}$ $m=\dfrac{63.1}{9\times 10^{16}}$ $m=7.01\times 10^{-16}$ kg 4) To convert from kg to protons: Mass of one proton $1.67\times 10^{-27}$ kg Number of protons $\dfrac{\text{total mass}}{\text{mass of one proton}} =$$\dfrac{7.01\times 10^{-16}}{1.67\times 10^{-27}} =$$4.20\times 10^{11}$ 5) Since the answer needs to be in terms of $x\times 10^{9}$: $4.20\times 10^{11}=420\times 10^{9}$ Therefore, $420\times 10^{9}$ protons would need to be converted to energy to release $63.1$ J of energy. The answer is 420. Inferred Instruction Calculate the number of protons (in multiples of $10^{9}$) required to release $63.1$ Joules of energy using the mass–energy equivalence equation and the mass of a proton. F1(Instruction, Inferred Instruction) 0.419 Rejected Response To solve this problem, we need to use the Einstein’s mass–energy equivalence formula: $E=mc^{2}$, where $E$ is the energy, $m$ is the mass, and $c$ is the speed of light. We are given the energy $E=63.1~\text{J}$. We need to find the mass $m$ that would release this energy when converted into energy using Einstein’s formula. First, we need to convert the speed of light c into meters per second. The speed of light is approximately $c=3\times 10^{8}~\text{m/s}.$ Now, we can rearrange the formula $E=mc^{2}$ to solve for mass $m$: $m=\frac{E}{c^{2}}.$ Plugging the given energy $E=63.1~\text{J}$ and the speed of light $c=3\times 10^{8}~\text{m/s}$ into the formula, we get: $m=\frac{63.1~\text{J}}{9\times 10^{16}{m^{2}/s^{2}}}=7.011\times 10^{-16}~\text{kg}.$ Now, we know that 1 proton has a mass of approximately $1.67\times 10^{-27}~\text{kg}.$ To find out how many protons have a mass of $7.011\times 10^{-16}~\text{kg}$, we can divide the total mass by the mass of one proton: Number of protons = $\frac{7.011\times 10^{-16}~\text{kg}}{1.67\times 10^{-27}~\text{kg/proton}}=4.20\times 10^{11}~\text{protons}.$ So, it would be necessary to convert the mass of $\boxed{4.20\times 10^{11}}$ protons into energy in order to release $63.1~\text{J}$ of energy. The answer is: $4.20\times 10^{11}$. Inferred Instruction Given an energy value (63.1 J), find out how many protons have a combined mass that, when converted to energy, would result in the given energy using Einstein’s mass-energy equivalence formula. F1(Instruction, Inferred Instruction) 0.375
Table 19: Example of FLIP on the RewardBench2 Math subset. The chosen response correctly converts the final answer into multiples of $10^{9}$, whereas the rejected response does not. FLIP successfully captures this difference, yielding higher similarity between the inferred instruction based on the chosen response and the original instruction than that based on the rejected response.